OpenStudy (anonymous):

If P(a|b,c) = P(a) then prove that P(b|c) = P(b). Can anyone give a step by step explanation to prove this?

OpenStudy (zarkon):

Not true

OpenStudy (anonymous):

Ok..could you please elaborate your answer?

OpenStudy (anonymous):

Or better if you can give some counter example..I would really appreciate that.

OpenStudy (zarkon):

roll a die and flip a coin let A={roll a 2} B=head C=tail then P(A|B,C)=P(A) since A is independent of B and C but P(B|C)=0\(\neq\frac{1}{2}\)=P(B)